OpenStudy (anonymous):

A car travels 100 meters in 5 seconds. What is the speed of the car? speed = distance over time; 1 mile = 1609 meters 8.9 miles per hour 16 miles per hour 20 miles per hour 44.7 miles per hour

OpenStudy (anonymous):

do you know the equation for linear speed?

OpenStudy (anonymous):

sorry, of course you do silly me

OpenStudy (anonymous):

first find out how many meters that is

OpenStudy (nurali):

So how much is 100m in miles?? 100/1609 right? :D And how much is 5 seconds in hours? 5/3600 right? :D So miles/hour = \[\frac{ 100 }{ 1609 }*\frac{ 3600 }{ 5 }\] 44.7 miles

OpenStudy (anonymous):

what naruli said is the better answer

OpenStudy (anonymous):

I learned it wrong, and I still have to remember when it is a good time to use fractions or decimals. Fractions are better here.